### v1

# Evaluating and Improving Automatic Speech Recognition Systems for Korean Meteorological Experts

Abstract
This paper explores integrating Automatic Speech Recognition (ASR) into natural language query systems to improve weather forecasting efficiency for Korean meteorologists. We address challenges in developing ASR systems for the Korean weather domain, specifically specialized vocabulary and Korean linguistic intricacies. To tackle these issues, we constructed an evaluation dataset of spoken queries recorded by native Korean speakers. Using this dataset, we assessed various configurations of a multilingual ASR model family, identifying performance limitations related to domain-specific terminology. We then implemented a simple text-to-speech-based data augmentation method, which improved the recognition of specialized terms while maintaining general-domain performance. Our contributions include creating a domain-specific dataset, comprehensive ASR model evaluations, and an effective augmentation technique. We believe our work provides a foundation for future advancements in ASR for the Korean weather forecasting domain. [^1] Our code and dataset will be publicly available.
[…]
Evaluating and Improving Automatic Speech Recognition Systems
for Korean Meteorological Experts
ChaeHun Park       Hojun Cho       Jaegul Choo KAIST AI {ddehun,hojun.cho,jchoo}@kaist.ac.kr
1 Introduction
Meteorological experts utilize diverse data sources for weather forecasting. The database systems they use are vast and complex, making it challenging for those without extensive system knowledge to access the necessary information. Natural language query systems can assist the experts by allowing them to input queries in natural language. The system translates these queries into database queries (e.g., SQL) to search the database (Zhong et al., 2017; Kim et al., 2020; Deng et al., 2022). This process enables forecasters to efficiently access and interpret complex databases without needing to write complex and structured queries themselves. Notably, Jo et al. (2023) developed an integrated search system for Korean weather data, enabling users to query extensive meteorological information using natural language. This system simplifies data retrieval and thereby improves operational efficiency.
Figure 1: Qualitative comparison of different models for Korean meteorological queries. All models are based on Whisper large-v2 (Radford et al., 2023). The models compared include Zero-shot, fine-tuned on a General domain dataset (i.e., KsponSpeech (Bang et al., 2020)), and fine-tuned on a Weather domain dataset synthetically generated by a text-to-speech API. Wrongly predicted words are manually highlighted in strikethrough by the authors.
Incorporating Automatic Speech Recognition (ASR) into these systems can improve efficiency by enabling voice-activated queries, which is beneficial for meteorologists who often multitask. However, developing an effective ASR system for the Korean weather domain presents challenges. Most existing ASR models, primarily trained on general domains, may struggle with the specialized vocabulary and pronunciation of meteorological terms (Yadav and Sitaram, 2022). Additionally, despite being trained in multiple languages, many recent multilingual ASR models often underperform in non-English languages, including Korean, due to their unique linguistic characteristics and limited representation in training data (Li et al., 2021; Radford et al., 2023; Ferraz et al., 2024; Song et al., 2024). The agglutinative nature and phonetic complexity of Korean add to these challenges (Lee et al., 2019; Cho et al., 2020). Therefore, the combination of Korean language characteristics and domain-specific vocabulary in meteorology makes developing an effective ASR system particularly difficult.
To investigate the extent of these challenges, we constructed a specialized evaluation dataset to assess how well ASR models perform in the Korean weather domain. We recorded 2,000 spoken queries derived from the natural language dataset created by Jo et al. (2023). Five native Korean speakers recorded the queries, following pronunciation guidelines from domain experts to ensure accurate pronunciation of domain-specific terms. Our evaluation dataset underwent rigorous manual validation to ensure its accuracy and reliability, making it suitable for evaluating ASR systems specifically in the meteorological domain.
Using this dataset, we evaluated popular multilingual ASR models (Radford et al., 2023) in different configurations. Training ASR models with general-domain Korean speech data (Bang et al., 2020) led to decreased performance, highlighting the challenges posed by distributional shifts in specialized domains. For instance, as shown in Fig. 1, both a zero-shot model and a model fine-tuned on a general domain dataset failed to accurately transcribe domain-specific terms (e.g., weather map, average wind speed, and KIM (Korean Integrated Model)) and unit system (e.g., hPa). These results revealed performance issues with domain-specific terminology, underscoring the need for models tailored to this specialized vocabulary. To resolve this issue, we employed a text-to-speech-based data augmentation method, which improved the recognition of specialized terms while maintaining performance on general-domain ASR tasks. We believe our findings provide valuable insights for developing effective domain-specific ASR systems, particularly within the Korean weather forecasting domain. Our contributions can be summarized as follows:
  • We developed a specialized evaluation dataset for Korean weather-related ASR tasks by recording and validating spoken queries from native Korean speakers.
  • We conducted evaluations using different model configurations, identifying areas where performance could be improved with domain-specific dataset creation.
  • We proposed and demonstrated the effectiveness of a text-to-speech-based data augmentation method to improve ASR performance in recognizing specialized meteorological terms.
2 Related Work
Speech Recognition for Specialized Domains
Speech recognition technology has been extensively studied in various specialized domains, such as medical (Le-Duc, 2024) and financial fields (O’Neill et al., 2021). One major challenge is accurately handling domain-specific terminology. Zheng et al. (2021) demonstrated that synthetic speech data enhances the recognition of out-of-vocabulary words in domain-specific contexts. Shamsian et al. (2024) propose a keyword-guided adaptation to improve the recognition accuracy of models for specialized terms. Our research specifically focuses on the Korean meteorological domain, creating an evaluation dataset tailored to assess ASR performance in this specialized field.
Speech Recognition for Korean
Research on ASR for non-English languages, including Korean, presents unique challenges. Multilingual ASR models, such as Whisper (Radford et al., 2023), often underperform with languages like Korean due to unique linguistic characteristics and limited representation in training datasets (Li et al., 2021; Yadav and Sitaram, 2022; Ferraz et al., 2024; Song et al., 2024). This performance drop is often attributed to the lower proportion and diversity of non-English data in training sets. Efforts to improve ASR for Korean include datasets like KsponSpeech, a spontaneous speech corpus, and ClovaCall, a goal-oriented dialog speech corpus, provide valuable resources for developing and evaluating Korean ASR systems (Bang et al., 2020; Ha et al., 2020). Our research evaluates ASR performance specifically within the Korean meteorological domain, using a tailored evaluation dataset to identify and address domain-specific performance limitations.
Figure 2: A construction pipeline of ASR evaluation dataset for Korean meteorological domain.
3 Dataset Construction
This study aims to develop a specialized dataset for evaluating Automatic Speech Recognition (ASR) systems within the Korean meteorological domain. The dataset comprises 2,000 spoken queries derived from natural language questions. The construction process includes several key steps: utilizing the natural language question dataset from Jo et al. (2023), obtaining pronunciation guidelines for specialized terms from domain experts, collecting speech data from native speakers, and verifying the correctness of the recorded audio. The detailed data construction pipeline (section 3.1) and the analysis of the resulting dataset (section 3.2) are as follows.
3.1 Data Construction Pipeline
Acquiring Natural Language Questions
We began by utilizing the natural language question dataset released by Jo et al. (2023). This dataset includes a diverse range of questions relevant to the Korean meteorological domain, providing a robust foundation for our spoken query dataset. The dataset consists of two subsets: URL and SQL. We selected 1,300 questions from the URL subset and 700 questions from the SQL subset to ensure comprehensive coverage of various meteorological topics that forecasters typically seek. Note that the original dataset in Jo et al. (2023) contains natural language queries mapped to their corresponding structured queries (e.g., SQL). For our purposes, we used only natural language queries, focusing on how these questions would be articulated by users in real-world scenarios. For more details about natural language questions, please refer to the original paper of the dataset (Jo et al., 2023).
Obtaining Pronunciation Guidelines from Domain Experts
To ensure the accurate pronunciation of specialized meteorological terminology, we collaborated with domain experts. These experts provided comprehensive guidelines on how to pronounce specific terms correctly. For example, the acronym "BUFR" should be pronounced as "buffer" and "AMI" as "A-M-I." Additionally, terms such as "33009 station" could be pronounced as "three three zero zero nine station" or "thirty-three thousand and nine station." These detailed instructions were crucial for maintaining the linguistic integrity of the dataset, ensuring that all domain-specific terms were articulated accurately by the speakers. More detailed guidelines are provided in Appendix A.
Uttr. time (sec) 7.212.55 - min time 2.12 - max time 19.18 Avg. words 8.382.97 Avg. chars 32.0012.50 Unique words 2759 Absent ratio (%) 34.43
Table 1: Dataset Statistics. Uttr. time (sec) denotes the audio lengths. Words are tokenized using the NLTK library (Bird and Loper, 2004). Absent ratio (%) refers to the percentage of unique words in our dataset that are not present in the Korean ASR dataset for open-domain dialog (Bang et al., 2020), divided by the total number of unique words in our dataset. Numbers in subscript represent standard deviation values.
Collecting Speech Data from Native Speakers
Five native Korean speakers were recruited to record the spoken queries. Four were male, and one was female, all in their twenties. Each participant was given the pronunciation guidelines and detailed instructions on how to record the questions. Three participants recorded 500 queries each, while the other two recorded 200 and 300 queries, respectively. Participants were compensated at a rate of 16,000 KRW (approximately 11.6 USD) per 100 queries. The recordings were conducted in controlled environments to minimize background noise and ensure high audio quality. This phase resulted in a substantial collection of audio recordings that correspond to the natural language questions.
Verifying the Recorded Voices
Following the collection of speech data, a rigorous verification process was implemented. Human verifiers listened to each recording to ensure accuracy and adherence to the pronunciation guidelines. They also checked for clarity and the absence of background noise or errors. Two of the authors are employed for this verification process. This process was essential to guarantee the quality and reliability of the dataset for subsequent ASR evaluation.
3.2 Dataset Analysis
Table 1 presents the overall statistics of our dataset, which consists of 2,000 spoken queries. The average utterance time is 7.21 seconds, ranging from 2.12 to 19.18 seconds, with each utterance containing an average of 8.38 words and 32.00 characters. The proportion of unique words in our dataset that are not found in the general domain ASR dataset (Bang et al., 2020) is 34.43%, highlighting the specialized and domain-specific nature of our dataset.
Table 2 shows the top 10 most frequently occurring words in both general and weather domain datasets. The general domain contains common conversational words, whereas the weather domain includes specialized terminology, underscoring the necessity of domain-specific datasets for improving ASR performance in meteorological applications.
General thing (거)    but (근데)    not (안)
that (그)    just (그냥)    I (내가)
such (그런)    uh (어)    what (뭐)
Weather surface (지상)   Standard Normals (평년값)
ocean (해양)   before (이전)   after (이후)
radar (레이더)   weather chart (일기도)
inform (알려줘)   satellite (위성)
Table 2: Ten most frequently occurred words in Korean ASR datasets in open-domain dialog (General) and weather (Weather) domains. The KsponSpeech (Bang et al., 2020) is used as a general domain dataset. Each English word is manually translated by the authors from the original Korean word presented in the bracket.
4 Experimental Setup
4.1 Evaluation Datasets
We use our Korean ASR dataset for the weather domain to evaluate different ASR models. The dataset comprises 2,000 spoken queries, which we randomly split into 500 samples for the development set and 1,500 samples for the test set. Additionally, we utilize the eval-clean test set from the KsponSpeech dataset (Bang et al., 2020) to evaluate the models in a general conversation domain. This dataset consists of 3,000 audio and their corresponding answer transcriptions. This approach allows us to compare models across both domain-specific and general ASR tasks.
4.2 Metrics
We use Character Error Rate (CER) and Word Error Rate (WER) as evaluation metrics to measure the distances between the prediction and ground-truth transcription. Additionally, we employ a space-normalized Word Error Rate (sWER) (Bang et al., 2020), which accounts for the flexibility and variations of space rules in Korean.
4.3 Models and Training Datasets
Models
For our experiments, we utilized the multilingual Whisper model family (Radford et al., 2023), specifically leveraging four different model sizes: tiny (39M), small (244M), medium (769M), and large-v2 (1550M). These pre-trained models were evaluated in a zero-shot manner to assess their initial performance without additional training for our dataset. The target language and task in the prefix tokens of Whisper are set to Korean and transcribe, respectively.
Fine-tuning on General Open-domain Dialogues
Besides zero-shot evaluation setup, to explore the impact of adapting a multilingual ASR model to the Korean language, we fine-tuned an ASR model using the KsponSpeech (Bang et al., 2020) dataset. The dataset contains 619k training instances about open-domain dialogue utterances from Native Korean annotators. The Whisper large-v2 is fine-tuned on this dataset.
Model Params. KsponSpeechEval (3k) WeatherDev (0.5k) WeatherTest (1.5k)
CER WER sWER CER WER sWER CER WER sWER
Tiny 39M 32.94 55.26 51.59 35.12 78.74 58.14 34.06 77.61 56.95
Small 244M 16.88 33.56 27.51 16.92 47.45 29.39 17.00 48.56 29.61
Medium 769M 14.83 31.12 23.08 13.21 40.77 20.82 13.37 41.72 21.10
Large-v2 1550M 15.56 31.13 22.78 11.83 37.00 19.20 11.99 38.38 18.41
Table 3: Zero-shot ASR evaluation results of the Whisper model family. We used CER, WER, and sWER as evaluation metrics, where lower scores indicate better performance. The Weather dataset denotes our Korean weather-domain evaluation dataset. Whisper models of different sizes were evaluated. The lowest and the second-lowest scores are highlighted in bold and underlined, respectively.
Figure 3: Zero-shot qualitative results of pre-trained Whisper models with different model sizes. The answer queries in English are manually translated from the original Korean answer. Incorrect model predictions are highlighted with strikethrough. Omitted correct answers are highlighted with underline. Spacing errors are not highlighted for better visualization. Best viewed in color.
Data Augmentation for Meteorological Domain
Inspired by Zheng et al. (2021), we hypothesize that teaching the model the pronunciation of specialized terms in the weather domain is crucial for accurate transcription. To facilitate this, we used a simple text-to-speech (TTS) system to generate audio for natural language queries in the weather domain, which was then used for training. We selected 10k and 9.8k instances from the URL and SQL subset of Jo et al. (2023) as the training queries. To avoid test set leakage, the natural language queries in our human-recorded evaluation dataset were excluded from these training queries. Using the Google TTS service [^2] https://gtts.readthedocs.io/, we generated audio for the selected queries. English words in the queries were converted to their Korean pronunciations before being provided to the TTS service based on expert guidelines described in Section 3.1. The resulting TTS-generated audio dataset was used as a training dataset. We fine-tuned the Whisper large-v2 either solely on the TTS-based datasets or on both the TTS-based datasets and the KsponSpeech dataset, resulting in two different ASR models.
4.4 Implementation Details
All fine-tuned models are parameter-efficiently trained using LoRA (Hu et al., 2022) with $r$=32. The fine-tuning process involved 3 epochs of training with a batch size of 48, a learning rate of 1e-3, and a warmup ratio of 0.1 with AdamW optimizer (Loshchilov and Hutter, 2017). For fine-tuned models, we saved the checkpoint after each epoch and selected the best one based on the lowest CER score from the development set of our dataset. The greedy decoding (Holtzman et al., 2020) is used as a decoding algorithm for all models. All models are implemented with Transformers framework (Wolf et al., 2020) and PyTorch (Paszke et al., 2019).
Figure 4: Qualitative results of ASR models with different fine-tuning datasets. The answer queries in English are manually translated from the original Korean answer. ZS denotes the pre-trained Whisper large-v2 model in zero-shot evaluation. FT (General) and FT (Both) denote Whisper large-v2 models fine-tuned on KsponSpeech and both KsponSpeech and TTS-based Weather datasets, respectively. The indicators are the same as Fig. 3.
5 Results
Main Results
Table 3 shows the zero-shot evaluation results of Whisper models with different sizes in general and weather domains. We observed that model performance improved with increasing model size across both datasets. For example, in the KsponSpeech dataset, the Whisper large-v2 model achieved a CER of 15.56, compared to 32.94 for the Tiny model. Table 4 illustrates the impacts of fine-tuning datasets. Fine-tuning the Whisper large-v2 model on the KsponSpeech dataset significantly improved performance in the open-domain dialogue domain, reducing the CER to 10.48. However, this led to a decrease in performance in the weather domain, with a CER of 16.83. In contrast, fine-tuning on the TTS-based Weather dataset enhanced performance in the weather domain, achieving a CER of 6.56, while slightly reducing performance in the general domain. Combining fine-tuning on both KsponSpeech and the TTS-based Weather dataset provided the best overall results, showing substantial improvements over zero-shot performance in both domains, with a CER of 8.28 in the weather domain and 10.54 in the general domain. These findings suggest that increasing model size and using a combination of general and domain-specific training data effectively enhance ASR performance, offering meaningful gains in both general and specialized applications.
Fine-tune Data KsponSpeechEval WeatherTest General Weather CER WER CER WER 15.56 31.13 11.99 38.38 10.48 24.05 16.83 51.50 15.36 32.26 6.56 18.69 10.54 23.94 8.28 21.78
Table 4: Impacts of fine-tuning datasets. The Whisper-large-v2 model is used for the experiments. Models fine-tuned on Ksponspeech and TTS-based Weather datasets are marked with ✓in General and Weather, respectively. The indicators are the same as Table 3.
Qualitative Results
The qualitative results presented in Fig. 3 and Fig. 4 highlight the challenges faced by ASR models in accurately recognizing domain-specific terms. Fig. 3 shows results from pre-trained Whisper models of different sizes, while Fig. 4 includes models fine-tuned on the general domain (KsponSpeech) and both general and weather domains (KsponSpeech + TTS-based weather dataset). Zero-shot models often struggled with domain-specific terminology, frequently predicting incorrect but similar-sounding words. For instance, in the third example of Fig 3, the Tiny model mispredicted "입자계수기" ("ip-ja-gye-su-gi", "particle counter") as "입장해두기" ("ip-jang-hae-du-gi", admission). In the first example, the Large model incorrectly split the place name "백령도" ("Baeng-nyeong-do", "Baengnyeong island") into "100" ("baek", "one hundred") and "0도" ("0 do", "0 degrees"). In Figure 4, the fine-tuned model using both general and weather datasets showed improved accuracy in recognizing specialized terms like "낙뢰" (nak-roe, "lightning") and "하순" ("ha-sun", "late order"), demonstrating the effectiveness of domain-specific data augmentation in enhancing ASR performance.
6 Conclusion
In this study, we developed a specialized evaluation dataset to assess ASR models for Korean meteorological forecasters’ natural language queries. Our experiments demonstrated that pre-trained models often struggle to recognize domain-specific meteorological terms, highlighting the need for efforts to build reliable ASR models in specialized applications. The use of synthetic training samples for the weather domain showed some effectiveness in improving the accuracy of these specialized terms. These findings emphasize the importance of tailored evaluation datasets and domain-specific adaptations in developing robust ASR systems to support meteorologists in their forecasting tasks.
Appendix A Pronunciation Guideline Details for Speech Data Collection
To ensure accurate pronunciation of meteorological terms, we consulted domain experts from the National Institute of Meteorological Sciences (NIMS). The focus was primarily on English words and abbreviations, along with various unit and number expressions. We note that a single word or expression can be pronounced in different ways. Experts provided detailed pronunciation guidelines for 256 words and expressions frequently used in the Korean weather domain. These guidelines were given to annotators as an initial reference and were available throughout the recording process. Selected examples from the pronunciation guidelines are shown in Fig. 5.
Figure 5: Samples of pronunciation rules for meteorological terminologies. The Pronunciation (English) are manually written by the authors for clarity and were not provided at the data annotation phase.